# New Tight Bounds for SGD without Variance Assumption: A Computer-Aided Lyapunov Analysis

### v1

@@ D.1 Bounds for any step-size
  • $a_{t}=\phi^{2(T-t)}$, where $\phi^{2}=\phi_{opt}^{2}+\varepsilon$, with
    $\phi_{opt}=\max\{1-\gamma\mu;\gamma L-1\}\quad\text{ and }\quad\varepsilon\in[0,1-\phi^{2}_{opt}),$
@@
    $\alpha\geq\max\left\{\frac{\gamma^{2}}{2}\left(1+\frac{\phi_{opt}^{2}}{\varepsilon}\right),\frac{1-\phi^{2}_{opt}-\varepsilon}{2L^{2}}\right\},$
@@
Then the sufficient Lyapunov conditions of Theorem B.4. Moreover, we have
$\phi\xrightarrow{\varepsilon\to 0}\phi_{opt},\quad\alpha\xrightarrow{\varepsilon\to 0}+\infty,\quad e\xrightarrow{\varepsilon\to 0}\gamma^{2}.$
Proof.
The arguments in this proof are mostly recycled from Proposition D.1. In particular, we shall use the same values for $\rho$ and $a_{t}$. Regarding the value of $\alpha$, we impose $a\alpha^{2}-2b\alpha+c\leq 0$ where $a,b,c$ are defined in Equation (7), which we recall for convenience:
$\begin{cases}a&=(L-\mu)^{2},\\ b&=\gamma(L+\mu)-\mu L\gamma^{2}-(1-\phi^{2}),\\ c&=\gamma^{2}(\phi_{opt}^{2}+\varepsilon).\end{cases}$
@@
$b=2\gamma L-L^{2}\gamma^{2}-(1-\phi_{opt}^{2})+\varepsilon=2\gamma L-L^{2}\gamma^{2}-\gamma L(2-\gamma L)+\varepsilon=\varepsilon.$
@@
$\alpha\geq\frac{c}{2b}=\frac{\gamma^{2}(\phi_{opt}^{2}+\varepsilon)}{2\varepsilon}=\frac{\gamma^{2}}{2}\left(1+\frac{\phi_{opt}^{2}}{\varepsilon}\right).$ (8)
It then remains to verify the conditions of Theorem B.4. Conditions 1 and 2 are readily satisfied. Condition 3 can be rewritten as
$\alpha\geq\frac{1-\phi^{2}_{opt}-\varepsilon}{2L^{2}},$
thus justified the bound on $\alpha$. Condition 4 requires that $\alpha\geq\tfrac{\gamma^{2}}{2}$, which is already true because of Equation (8). Condition 5 is satisfied under Equation (8). Finally, Condition 6 is satisfied with
@@
where the denominator is strictly positive because of (8). ∎
Theorem D.3 (Bound for SGD. Strongly convex case, general step-sizes).
Let Assumptions A.5 and A.7 hold, with $\mu>0$. Let $\varepsilon\geq 0$ and $\gamma_{opt}=\tfrac{2}{\mu+L}$. Let $x_{t}$ be generated by SGD, with $\gamma L\in(0,2)$. Assume that either $\varepsilon>0$ or $\gamma\neq\gamma_{opt}$. Then, for every $T\geq 1$,
@@
where $\phi^{2}=\varepsilon+(\max\{1-\gamma\mu;\gamma L-1\})^{2}\in[0,1)$, and $e$ is defined in Proposition D.1 and verifies
$e\leq\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)^{2}}{4\varepsilon+(L+\mu)(L-\mu)\gamma|\gamma-\gamma_{opt}|}\right).$
Proof.
@@
  • Consider $L>\mu$. Proposition D.1, combined with Theorem B.4, provides parameters ensuring that the Lyapunov energy decreases with, in particular, $a_{T}=1$, $a_{0}=\phi^{2T}$. We deduce the wanted bound from Lemma B.3.
  • Consider $L=\mu$. Let us consider first $\varepsilon>0$. Combining Lemma B.3, Theorem B.4 and Proposition D.2 we obtain
@@
    where $e_{\varepsilon}$ converges to $\gamma^{2}$ and $\phi\to\phi_{\text{opt}}$ when $\varepsilon\to 0$, thus yielding the wanted bound when $\varepsilon=0$. If we wish to prove the claim for $\varepsilon>0$, it is enough to realize that $\phi_{opt}\leq\phi$, together with the fact that $\tfrac{1-\phi^{2T}}{1-\phi^{2}}$ is nondecreasing with $\phi$.
The above propositions allow us to formulate bounds for SGD. As in the convex case, we consider optimal and non-optimal step-sizes as separate cases.
Theorem D.4 (Tight bound for SGD. Strongly convex case, non-optimal step-sizes).
Let Assumptions A.5 and A.7 hold, with $\mu>0$. Let $x_{t}$ be generated by SGD, with $\gamma L\in(0,2)$ such that $\gamma\neq\gamma_{opt}=\tfrac{2}{\mu+L}$. Then, for every $T\geq 1$,
where $\phi=\max\{1-\gamma\mu;\gamma L-1\}\in[0,1)$, and
$e=\frac{2\gamma^{2}\phi}{|2-\gamma(L+\mu)|}=\gamma^{2}\left(1+\frac{\gamma(L-\mu)}{|\gamma-\gamma_{opt}|(L+\mu)}\right).$
Proof.
This is an immediate consequence of Theorem D.3 with $\varepsilon=0$. ∎
If $\gamma L\leq 1$, the above result can be simplified, allowing for an easier comparison with the literature.
Corollary D.5 (Tight bound for SGD. Strongly convex case, short step-size $\gamma L\leq 1$).
Let Assumptions A.5 and A.7 hold, with $\mu>0$. Let $x_{t}$ be generated by SGD, with $\gamma L\in(0,1]$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq(1-\gamma\mu)^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{2\gamma}{\mu(2-\frac{\mu}{L})}\,\sigma_{*}^{2}.$
Proof.
If $\gamma L\leq 1$, then $\phi=1-\gamma\mu$. Moreover,
$\frac{1-\phi^{2T}}{1-\phi^{2}}e\leq\frac{e}{1-\phi^{2}}=\frac{2\gamma^{2}(1-\gamma\mu)}{\gamma\mu(2-\gamma\mu)\big{(}2-\gamma(L+\mu)\big{)}}=\frac{2\gamma}{\mu}\left[\frac{(1-\gamma\mu)}{(2-\gamma\mu)\big{(}2-\gamma(L+\mu)\big{)}}\right].$
The rational function between the large brackets increases with $\gamma$ on $(0,\frac{1}{L}]$ (on $(0,\frac{2}{L+\mu})$, actually), so it can be bounded from above there by its value on $\gamma=\frac{1}{L}$, which is $\frac{1}{2-\frac{\mu}{L}}$. ∎
Finally, we provide a bound which remains valid for every steps-size, with the counterpart that the bias term gets sub-optimal.
Theorem D.6 (Subtight bound for SGD. Strongly convex case, general step-sizes).
Let Assumptions A.5 and A.7 hold, with $\mu>0$. Let $\varepsilon>0$, and let $x_{t}$ be generated by SGD, with $\gamma L\in(0,2)$. Then, for every $T\geq 1$,
where $\phi^{2}=\varepsilon+(\max\{1-\gamma\mu;\gamma L-1\})^{2}\in[0,1)$, and $e$ was defined in Proposition D.1 and verifies
$e\leq\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)^{2}}{4\varepsilon}\right).$
Proof.
This is an immediate consequence of Theorem D.3, where we used $\varepsilon>0$ to lower bound the denominator of $e$. ∎
D.2 Detailed comparison to previous works
Comparison to bach_non-asymptotic_2011.
In (bach_non-asymptotic_2011, Theorem 1), the authors analyze the convergence of SGD for strongly convex objectives without variance assumptions, aligning with the setting of our Theorem 3.3. They present the following bound:
$\mathbb{E}\mathopen{}\left[{\|x_{N}-x_{*}\|^{2}}\right]\mathclose{}\leq 2\exp(4L^{2}\gamma^{2}(N-1))\cdot\exp(-\mu\gamma N/4)\cdot(\|x_{0}-x_{*}\|^{2}+\sigma_{*}^{2}/L^{2})+4\gamma\sigma_{*}^{2}/\mu.$
Their proof is based on a Lyapunov energy function similar to the one employed in our analysis, though with different coefficients. The exponential form is presented for clarity in the context of their use of decreasing step-sizes. For the constant step-size scenario, their analysis yields
$\mathbb{E}\mathopen{}\left[{\|x_{N}-x_{*}\|^{2}}\right]\mathclose{}\leq\varphi^{N}\cdot\|x_{0}-x_{*}\|^{2}+2\gamma^{2}\cdot\frac{1-\varphi^{N}}{1-\varphi}\cdot\sigma_{*}^{2},$
where $\varphi=1-2\mu\gamma+2L^{2}\gamma^{2}$. It may be verified that $\phi^{2}\leq\varphi$, where $\phi$ is defined as in Theorem D.4, which connects their result to ours. In specific, our analysis explicitly establishes convergence over a broader range of step-sizes $\gamma$ compared to the range considered in their work, and provides a sharper rate.
Comparison to needell_stochastic_2016.
In (needell_stochastic_2016, Theorem 2.1), the authors obtain, under the same assumptions and for $\gamma<\frac{1}{L}$, the bound
$\mathbb{E}\mathopen{}\left[{\|x_{k}-x_{*}\|^{2}}\right]\mathclose{}\leq\left(1-2\gamma\mu(1-\gamma L)\right)^{k}\cdot\|x_{0}-x_{*}\|^{2}+\frac{\gamma}{\mu(1-\gamma L)}\cdot\sigma_{*}^{2}.$
We note that the bound is weaker than the one obtained in our results, and is restricted to short stepsizes.
Comparison to gower_sgd_2019.
The authors of (gower_sgd_2019, Theorem 3.1) study the quasi-strongly convex setting under the assumption of expected smoothness, which is tightly linked to our assumption (EC), as put forward in Remark A.6
For a constant step-size $\gamma\in(0,1/(2L)]$, they derive the bound
$\mathbb{E}\mathopen{}\left[{\|x_{N}-x_{*}\|^{2}}\right]\mathclose{}\leq(1-\gamma\mu)^{k}\|x_{0}-x_{*}\|^{2}+\frac{2\gamma}{\mu}\cdot\sigma_{*}^{2}.$
Our analysis extends the convergence result to a broader range of step-sizes. Furthermore, within this specific range of step-sizes, our derived convergence rate incorporates a squared term related to $(1-\gamma\mu)$, differing from the linear term presented in their bound. This is an effect often seen in simpler proofs.
Appendix E Performance Estimation Problem methodology
@@
In this section, we describe the Performance Estimation Problem methodology in more detail, and explain how we made use of it in this scenario. Most ideas are inspired by taylor_stochastic_2019.
@@
E.1 Problem reformulation
As we aim to obtain a computationally tractable problem, we do restrict ourselves to the finite-sum setting in this section. We do however note that the number of functions $m$ does not influence the final result, and can hence theoretically be taken to be infinite, thus recovering the results of Sections C and D.
As explained by Lemmas B.2 and B.3, our goal is to derive nonnegative parameters $\rho,(a_{t},e_{t})$ such that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\ldots,T-1$, where
@@ E.1 Problem reformulation
  • In the convex setting, Lemma B.2 tells us that we are interested in maximizing $\rho$, or equivalently minimizing $\tfrac{1}{\rho}$, where we fix $a_{0}=1$. In this setting we denote the bias and normalization by
@@
  • In the strongly convex setting, Lemma B.3 tells us we are interested in minimizing $a_{0}$, where we fix $a_{T}=1$. In this setting we denote the bias and normalization by
@@
$_{\text{opt}}=\text{Bias}_{\text{opt}}(T)\coloneq\inf_{\rho,(a_{t},e_{t})\in\mathcal{N}_{T}}\left\{\text{Bias}\right\}.$ Bias } .
@@
Note that $x_{0}$ is a variable of the follower problem, and hence acts as a slack variable for the constraint that $x_{t}$ is generated from $x_{0}$. Using this, and recalling the definition of $E_{t}$, we may write that
@@
for j{t,} and $\qquad\qquad\qquad\qquad\qquad\qquad\text{for $j\in\{t,*\}$ and $i=1,\ldots,m$},$
@@
The existence condition may seem like a complicated reformulation, but it allows us to use (taylor_smooth_2017, Theorem 4), a cornerstone of the performance estimation framework:
Theorem E.1.
@@
$\begin{dcases}\Delta_{t}&=\frac{a_{t+1}}{m}\sum_{i=1}^{m}(\bm{p}_{2m+1}-\gamma\bm{p}_{m+i})(\bm{p}_{2m+1}-\gamma\bm{p}_{m+i})^{T}-a_{t}\bm{p}_{2m+1}\bm{p}_{2m+1}^{T}\\ &\qquad\qquad\qquad-\frac{e_{t}}{m}\sum_{i=1}^{m}\bm{p}_{m+i}\bm{p}_{m+i}^{T},\\ \tilde{\Delta}&=\frac{\rho}{m}\sum_{i=1}^{m}\bm{f}_{i}-\bm{f}_{m+i},\\ A_{t,*}^{(i)}&=\bm{p}_{m+i}\bm{p}_{2m+1}^{T}+\frac{1}{2(L-\mu)}(\bm{p}_{i}-\bm{p}_{m+i})(\bm{p}_{i}-\bm{p}_{m+i})^{T}+\frac{\mu L}{2(L-\mu)}\bm{p}_{2m+1}\bm{p}_{2m+1}^{T}\\ &\qquad\qquad\qquad-\frac{\mu}{L-\mu}(\bm{p}_{i}-\bm{p}_{m+i})\bm{p}_{2m+1}^{T},\\ A_{*,t}^{(i)}&=-\bm{p}_{i}\bm{p}_{2m+1}^{T}+\frac{1}{2(L-\mu)}(\bm{p}_{i}-\bm{p}_{m+i})(\bm{p}_{i}-\bm{p}_{m+i})^{T}+\frac{\mu L}{2(L-\mu)}\bm{p}_{2m+1}\bm{p}_{2m+1}^{T}\\ &\qquad\qquad\qquad-\frac{\mu}{L-\mu}(\bm{p}_{i}-\bm{p}_{m+i})\bm{p}_{2m+1}^{T}.\\ \end{dcases}$
@@
$\tilde{B}_{t}=\quad\quad\quad\sup_{\Lambda_{t},\lambda_{t,*}^{(i)},\lambda_{*,t}}\quad$ 00 0 (Dual)
@@
We prove in Section E.3 that this is indeed the dual problem of Problem (Primal). We further prove that strong duality holds, meaning that $\tilde{B}_{t}=-B_{t}$. As a consequence,
@@
E.2 Converting PEP solutions into mathematical proofs

### v2

@@ A.2 Strongly Convex Setting
  • $\bullet$
    $a_{t}=\phi^{2(T-t)}$, where $\phi^{2}=\phi_{{\operatorname{opt}}}^{2}+\varepsilon$, with
    $\phi_{{\operatorname{opt}}}=\max\{1-\gamma\mu;\gamma L-1\}\quad\text{ and }\quad\varepsilon\in[0,1-\phi^{2}_{{\operatorname{opt}}}),$
  • $\bullet$
@@
    $\alpha=\max\left\{\frac{\gamma^{2}}{2}\left(1+\frac{\phi_{{\operatorname{opt}}}^{2}}{\varepsilon}\right),\frac{1-\phi^{2}_{{\operatorname{opt}}}-\varepsilon}{2L^{2}}\right\},$
  • $\bullet$
@@
Then the sufficient Lyapunov conditions of Theorem 3.1 are verified. Moreover, we have
$\phi\xrightarrow{\varepsilon\to 0}\phi_{{\operatorname{opt}}},\quad\alpha\xrightarrow{\varepsilon\to 0}+\infty,\quad e\xrightarrow{\varepsilon\to 0}\gamma^{2}.$
Proof
The arguments in this proof are mostly recycled from Proposition 8. In particular, we shall use the same values for $\rho$ and $a_{t}$. Regarding the value of $\alpha$, we impose $a\alpha^{2}-2b\alpha+c\leq 0$ where $a,b,c$ are defined in Equation (7), which we recall for convenience:
$\begin{cases}a&=(L-\mu)^{2},\\ b&=\gamma(L+\mu)-\mu L\gamma^{2}-(1-\phi^{2}),\\ c&=\gamma^{2}(\phi_{{\operatorname{opt}}}^{2}+\varepsilon).\end{cases}$
@@
$b=2\gamma L-L^{2}\gamma^{2}-(1-\phi_{{\operatorname{opt}}}^{2})+\varepsilon=2\gamma L-L^{2}\gamma^{2}-\gamma L(2-\gamma L)+\varepsilon=\varepsilon.$
@@
$\alpha\geq\frac{c}{2b}=\frac{\gamma^{2}(\phi_{{\operatorname{opt}}}^{2}+\varepsilon)}{2\varepsilon}=\frac{\gamma^{2}}{2}\left(1+\frac{\phi_{{\operatorname{opt}}}^{2}}{\varepsilon}\right).$ (8)
It then remains to verify the conditions of Theorem 3.1. Conditions 1 and 2 are readily satisfied. Condition 3 can be rewritten as
$\alpha\geq\frac{1-\phi^{2}_{{\operatorname{opt}}}-\varepsilon}{2L^{2}},$
thus justifying our definition of $\alpha$. Condition 4 requires that $\alpha\geq\tfrac{\gamma^{2}}{2}$, which is already true because of Equation (8). Condition 5 is satisfied under Equation (8). Finally, Condition 6 is satisfied with
@@
where the denominator is strictly positive because of Equation (8).
Theorem A.2(Strongly convex case, general step-size)
Let Assumptions 1 and 2 hold, with $\mu>0$. Let $\varepsilon\geq 0$ and $\gamma_{\operatorname{crit}}=\tfrac{2}{\mu+L}$. Let $x_{t}$ be generated by (SGD), with $\gamma L\in(0,2)$. Assume that either $\varepsilon>0$ or $\gamma\neq\gamma_{\operatorname{crit}}$. Then, for every $T\geq 1$,
@@
where $\phi^{2}=\varepsilon+(\max\{1-\gamma\mu;\gamma L-1\})^{2}\in[0,1)$, and $e$ is defined in Proposition 8 and verifies
$e\leq\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)^{2}}{4\varepsilon+(L+\mu)(L-\mu)\gamma|\gamma-\gamma_{\operatorname{crit}}|}\right)=\gamma^{2}\frac{\gamma^{2}_{\operatorname{crit}}\varepsilon+\gamma\gamma_{\operatorname{crit}}\phi_{\gamma}\phi_{\operatorname{crit}}}{\gamma^{2}_{\operatorname{crit}}\varepsilon+\gamma\phi_{\operatorname{crit}}|\gamma-\gamma_{\operatorname{crit}}|}.$
In particular, the bounds of Theorems 5.1 and 5.2 hold true.
Proof
@@
  • $\bullet$
    Consider $L>\mu$. Proposition 8, combined with Theorem 3.1, provides parameters ensuring that the Lyapunov energy decreases with, in particular, $a_{T}=1$, $a_{0}=\phi^{2T}$. We deduce the wanted bound from Lemma 8.
  • $\bullet$
    Consider $L=\mu$. Let us consider first $\varepsilon>0$. Combining Lemma 8, Theorem 3.1 and Proposition 9 we obtain
@@
    where $e_{\varepsilon}$ converges to $\gamma^{2}$ and $\phi\to\phi_{\operatorname{opt}}$ when $\varepsilon\to 0$, thus yielding the wanted bound when $\varepsilon=0$. If we wish to prove the claim for $\varepsilon>0$, it is enough to realize that $\phi_{{\operatorname{opt}}}\leq\phi$, together with the fact that $\tfrac{1-\phi^{2T}}{1-\phi^{2}}$ is nondecreasing with $\phi$.
The bounds of Theorems 5.1 and 5.2 now follow by either setting $\varepsilon=0$ or using that $\gamma=\gamma_{\operatorname{crit}}$.
We end by showing the bias-optimality claim of Theorem 5.1.
Proposition 10(Bias-optimality. Strongly convex case)
The bias obtained in the bound of Theorem 5.1 is optimal, in the sense that there exists a convex and smooth problem for which $m=1$ and the inequality is satisfied with equality.
Proof
It suffices to consider a one-dimensional quadratic function given by $f(x)=\frac{\mu}{2}\|x\|^{2}$ if $\gamma\leq\frac{2}{L+\mu}$, and by $f(x)=\frac{L}{2}\|x\|^{2}$ otherwise, and any $x_{0}\in\mathbb{R}^{d}$. The iterates $(x_{t})$ of gradient descent are given by $x_{t}=(1-\gamma\mu)^{t}x_{0}$ or $x_{t}=(1-\gamma L)^{t}x_{0}$. Noting that $x_{*}=0$ in either case, we obtain that
$\|x_{T}-x_{*}\|^{2}=\begin{cases}(1-\gamma\mu)^{2T}\cdot\|x_{0}-x_{*}\|^{2}&\quad\text{if $\gamma\leq\frac{2}{L+\mu}$}\\ (1-\gamma L)^{2T}\cdot\|x_{0}-x_{*}\|^{2}&\quad\text{if $\gamma\geq\frac{2}{L+\mu}$}\end{cases}=\max(1-\gamma\mu,\gamma L-1)^{2T}\cdot\|x_{0}-x_{*}\|^{2},$
matching the bias term obtained in Theorem 5.1.
Appendix B Performance Estimation Problem Methodology
@@
In this section, we describe the Performance Estimation Problem (PEP) methodology in more detail, and explain how we made use of it in this scenario. The methodology was initially introduced by drori_performance_2014 , further improved by taylor_smooth_2017 , and adapted by taylor_stochastic_2019 in the context of stochastic algorithms, that we will follow closely.
@@
B.1 Problem Reformulation
As we aim to obtain a computationally tractable problem, we do restrict ourselves to the finite-sum setting in this section. We do however note that the number of functions $m$ does not influence the final result, and can hence theoretically be taken to be infinite, thus recovering the results of Sections A.1 and A.2. In the same spirit, we here assume $\mathcal{H}$ to be a finite-dimensional euclidean space, although the proofs of our results directly extend to any real Hilbert space.
As explained by Lemmas 7 and 8, our goal is to derive nonnegative parameters $\rho,(a_{t},e_{t})$ such that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\ldots,T-1$, where
@@ B.1 Problem Reformulation
  • $\bullet$
    In the convex setting, Lemma 7 tells us that we are interested in maximizing $\rho$, or equivalently minimizing $\tfrac{1}{\rho}$, where we fix $a_{0}=1$. In this setting we denote the bias and normalization by
@@
  • $\bullet$
    In the strongly convex setting, Lemma 8 tells us we are interested in minimizing $a_{0}$, where we fix $a_{T}=1$. In this setting we denote the bias and normalization by
@@
$\text{Bias}_{\text{opt}}=\text{Bias}_{\text{opt}}(T)\coloneq\inf_{\rho,(a_{t},e_{t})\in\mathcal{N}_{T}}\left\{\text{Bias}\right\}.$
@@
Note that the initial point $x_{0}$ is a decision variable of the inner (follower) problem. As a result, the constraint that $x_{t}$ must be generated by running SGD for $t$ steps from $x_{0}$ is not restrictive. Indeed, for any target point $x\in\mathbb{R}^{d}$, one can construct an initial point $x_{0}\in\mathbb{R}^{d}$ and functions $f_{1},\dots,f_{m}\in\mathcal{F}_{\mu,L}$ such that the SGD iterates satisfy $x_{t}=x$. Consequently, allowing $x_{0}$ to vary freely effectively makes the SGD-generation constraint slack. Using this, and recalling the definition of $E_{t}$, we may write that
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\text{for $j\in\{t,*\}$ and $i=1,\ldots,m$},$
@@
The existence condition may seem like a complicated reformulation, but it allows us to use (taylor_smooth_2017, , Theorem 4), a cornerstone of the performance estimation framework:
Theorem B.1
@@
[…]
@@
$\tilde{B}_{t}=\quad\quad\quad\sup_{\Lambda_{t},\lambda_{t,*}^{(i)},\lambda_{*,t}}\quad$ $0$ (Dual)
@@
We prove in Section B.3 that this is indeed the dual problem of Problem (Primal). We further prove that strong duality holds, meaning that $\tilde{B}_{t}=-B_{t}$. As a consequence,
@@
B.2 Obtaining Mathematical Proofs